Assume the time required to pass through security at a particular airport follows the continuous uniform distribution with aThe answers are in red. Please do part a-f, Thanks in advance :)
Answer #1

solution f(x) = bra (M)=bae actor (numitorm loo[a, b] a = 10 0.038 b=36 b) Mean = atb - 10+36 = [23] stondend dari ben 160 g4